# Infinity-MM: Scaling Multimodal Performance with Large-Scale and High-Quality Instruction Data

### v1

@@ 3.3 Data Processing
After collecting all the data, we proceeded with data processing. First, to facilitate large-scale training, we standardized the format of data from various sources. Then, to improve training efficiency and enhance model performance, we conducted a series of data-cleaning steps. Specifically, we removed duplicate Image-Text pairs and filtered out images with high similarity based on their pHash values. Additionally, we used Qwen2-VL-2B to calculate the loss for each sample and excluded the top 5% with the highest loss, as high loss in well-trained multimodal models often indicates noisy data or outliers. To prevent data contamination, we also deduplicated the training set against images found in the test set. We have included more detailed data information in Appendix C.
4 Architecture and Training
@@ 4.1 Model Architecture
Aquila-VL builds upon LLaVA-OneVision architecture(Li et al., 2024a), comprising a language tower, a vision tower, and a projector.
@@
  • Language Tower We chose Qwen-2.5 (Bai et al., 2023a) as the language tower for its outstanding performance among open-source models and its availability in various sizes.
@@
  • Vision Tower We utilized SigLIP (Zhai et al., 2023), with approximately 400 million parameters, as the vision tower to extract visual features from input images and videos.
@@
  • Projector We utilized a two-layer MLP (Liu et al., 2024b) with a GELU (Hendrycks and Gimpel, 2023) activation to project visual features into the word embedding space.
@@
Models Params (B) Average MMBenchV1.1test MMStar MMMUval MathVistatestmini HallusionBench AI2Dtest OCRBench MMVet
DeepSeek-VL-1.3B Lu et al. (2024) 2.0 39.6 63.8 39.9 33.8 29.8 27.6 51.5 413 29.2
MiniMonkey Huang et al. (2024) 2.2 52.7 68.9 48.1 35.7 45.3 30.9 73.7 794 39.8
MiniCPM-V-2 Yao et al. (2024) 2.8 47.9 65.8 39.1 38.2 39.8 36.1 62.9 605 41.0
PaliGemma-3B-mix-448 Beyer* et al. (2024) 2.9 46.5 65.6 48.3 34.9 28.7 32.2 68.3 614 33.1
Phi-3-Vision Abdin et al. (2024) 4.2 53.6 65.2 47.7 46.1 44.6 39.0 78.4 637 44.1
InternVL2-2B Chen et al. (2024b) 2.1 53.9 69.6 49.8 36.3 46.0 38.0 74.1 781 39.7
H20VL-Mississippi-2B Galib et al. (2024) 2.1 54.4 64.8 49.6 35.2 56.8 36.4 69.9 782 44.7
XinYuan-VL-2B Cylingo (2024) 2.1 56.1 75.4 51.9 43.6 47.1 36.0 74.2 782 42.7
Qwen2-VL-2B Wang et al. (2024a) 2.1 57.2 72.7 47.8 41.7 47.9 41.5 74.6 810 50.7
Aquila-VL-2B 2.1 59.5 75.2 54.9 47.4 59.0 43.0 75.0 772 44.3
Table 3: Performance comparison between Aquila-VL-2B and other models. The results are cited from the official leaderboad of VLMEvalKit and Galib et al. (2024).
@@
w/o Synthetic Data 57.1 71.2 53.0 44.2 58.4 37.9 74.3 773 40.0
Table 4: Comparison of the impact on model performance with and without synthetic data.
4.2 Training Details
We implemented a curriculum learning approach to train Aquila-VL-2B in phases following Wang et al. (2024a); Li et al. (2024b). Our training is divided into four stages, progressively increasing the task difficulty, image resolution, and data quality. The training setup is presented in Table 2.
  • Stage 2: We utilized general visual instruction data for further training to equip the model with fundamental capabilities for solving multimodal tasks. The data was divided into three subsets, and during each stage of training, the maximum visual resolution was progressively increased to enhance the model’s comprehension of visual information.
  • Stage 3: We employed selective visual instruction data for training and further increased the maximum resolution to improve performance.
  • Stage 4: We fine-tuned the model using training data from GPT-4 and synthetic data. Experiments demonstrate that scaling with synthetic data can further enhance model performance.
We used the official codes of LLaVA-OneVision [^1] https://github.com/LLaVA-VL/LLaVA-NeXT/tree/main/scripts/train. Besides, we also supported the training of Aquila-VL-2B in FlagScale BAAI (2024a), which is a comprehensive toolkit tailored for large models based on open-source projects. Under identical configurations, end-to-end training with FlagScale achieves a 1.7x acceleration compared to DeepSpeed. To efficiently handle large-scale training data without encountering out-of-memory (OOM) errors, we have developed an optimized multi-modal data loader, leveraging Megatron Energon Nvidia (2024). This loader processes datasets in a pre-formatted structure, prepared offline, to ensure efficient data consumption during training. The model was trained on both Nvidia A100 and chips from a Chinese manufacturer.
5 Evaluation
In this section, we first evaluate the performance of the model through a comparative analysis of multiple benchmarks, demonstrating the advantages of our approach. Subsequently, we conduct a detailed examination of the model’s specific capabilities, including general visual perception, document understanding and mathematical reasoning. Finally, we carry out an ablation study to investigate several key components of our approach.
5.1 Compare to SOTAs
We assessed the visual capabilities of Aquila-VL-2B using a range of visual benchmarks provided by the VLMEvalKit Duan et al. (2024). Aquila-VL-2B demonstrates highly competitive performance at the same scale, achieving new state-of-the-art results. Specifically, we evaluated the capabilities of Aquila-VL-2B across three task categories.
General Visual Question Answering We conducted extensive evaluations across a diverse array of general visual question answering benchmarks: MMStar, HallusionBench, MMVet Yu et al. (2023), and MMBench-1.1 Liu et al. (2023b). Aquila-VL-2B demonstrated strong performance across these benchmarks, achieving or surpassing state-of-the-art results in most cases at the same scale. On MMStar, which evaluates multimodal capabilities by integrating visual and textual information, Aquila-VL-2B achieved a remarkable score of 54.9, surpassing previous state-of-the-art results and demonstrating its strong proficiency in handling diverse multimodal tasks. On HallusionBench, which evaluates image-context reasoning, Aquila-VL-2B achieved a score of 43.0, surpassing both the previous state-of-the-art and strong baselines, demonstrating its superior ability in understanding and reasoning within complex visual contexts. On MMVet, which evaluates large multimodal models for integrated vision-language capabilities across 16 complex multimodal tasks, Aquila-VL-2B achieved a score of 44.3, demonstrating its ability to address diverse multimodal challenges. On MMBench, which evaluates fine-grained abilities across 20 dimensions, Aquila-VL-2B exhibited strong performance, achieving a score of 76.3 on the English test set, matching the state-of-the-art, and 74.1 on the Chinese test set, demonstrating its robust capabilities in this benchmark.
Knowledge and Mathematical Reasoning We conducted experiments on the AI2D Kembhavi et al. (2016), MMMU Yue et al. (2023), and MathVista datasets to evaluate the model’s capabilities in knowledge and mathematical reasoning. The MMMU dataset is a new benchmark designed to assess multimodal models on extensive multi-disciplinary tasks that require college-level subject knowledge and deliberate reasoning. Aquila-VL-2B achieved a strong score of 47.4, surpassing state-of-the-art results at the same scale and demonstrating its proficiency in addressing complex multimodal challenges. MathVista is a comprehensive benchmark comprising 6,141 diverse examples of mathematical and visual tasks. The Aquila-VL-2B series exhibited exceptional performance on the MathVista benchmark, achieving a score of 59.0, thereby outperforming other large vision language models (LVLMs). AI2D focuses on multiple-choice questions related to scientific diagrams containing text. Aquila-VL-2B exhibited outstanding performance at a comparable scale, achieving a score of 75.0, which represents the highest performance in this benchmark, highlighting its competitive strengths.
Text Reading We assessed Aquila-VL-2B’s capabilities in text reading and diagram comprehension using the OCRBench dataset. OCRBench is a mixed-task dataset that emphasizes mathematical formula parsing and information extraction, in addition to text-based visual question answering (VQA). The performance results highlight potential areas for further optimization.
5.2 Effects of Sythesis Data
@@
To assess the impact of synthetic data on model performance, we conducted an ablation study. In this experiment, we removed all synthetic data and trained the model using only the original GPT-generated data. The results, as shown in Table 4, revealed a significant decline in overall model performance after removing the synthetic data. This demonstrates that the synthetic data played a crucial role in enhancing the model’s performance, further validating the effectiveness of our approach in data augmentation and diversity.
@@ 5.3 Data Scaling
Figure 3: The change of model performance with training data size.
To further analyze the impact of data size scaling on model performance, we conducted a detailed study on how model performance varies with the amount of training data. The results, shown in Figure 3, indicate a consistent improvement in performance as the training data increases. This trend clearly demonstrates that expanding the scale of instruction data has a significant positive effect on model performance. This observation suggests that as more diverse instruction data is introduced, the model’s ability to handle complex tasks is enhanced. Therefore, scaling up the instruction data is an effective strategy for improving overall model performance.
@@ 6 Conclusion
In this work, to enhance the performance of open-source models, we built the Infinity-MM multimodal instruction dataset with tens of millions of samples, increasing the data volume to improve model efficacy. Besides, we proposed a method for synthesizing instruction data based on open-source models, which further generated high-quality instruction data and expanded the dataset size. Ultimately, we trained the Aquila-VL-2B model using Infinity-MM, achieving state-of-the-art performance for models of comparable size.
Appendix A Comprehensive Benchmark Comparisons to SOTAs
In Table 5, we provide a more comprehensive assessment of our model’s performance through extensive comparisons across multiple benchmarks against SOTA models. These benchmarks encompass a diverse range of tasks, including visual recognition, reasoning, and document understanding, offering a robust evaluation of the model’s capabilities.
Capability Benchmark MiniCPM-V-2 InternVL2-2B XinYuan-VL-2B Qwen2-VL-2B-Instruct Aquila-VL-2B
GeneralVQA MMBench-ENtest 69.4 73.4 78.9 74.9 78.8
MMBench-CNtest 65.9 70.9 76.1 73.9 76.4
MMBench_V1.1test 65.2 69.7 75.4 72.7 75.2
MMT-Benchtest 54.5 53.3 57.2 54.8 58.2
RealWorldQA 55.4 57.3 63.9 62.6 63.9
HallusionBench 36.8 38.1 36.0 41.5 43.0
SEEDBench2Plus 51.8 60.0 63.0 62.4 63.0
LLaVABench 66.1 64.8 42.4 52.5 68.4
MMStar 41.6 50.2 51.9 47.8 54.9
POPE 86.6 85.3 89.4 88.0 83.6
MMVet 44.0 41.1 42.7 50.7 44.3
Knowledge& Mathematical MMMUval 39.6 34.9 43.6 41.9 47.4
ScienceQAtest 80.4 94.1 86.6 78.1 95.2
AI2Dtest 64.8 74.4 74.2 74.6 75.0
MathVistatestmini 39.0 45.0 47.1 47.9 59.0
MathVersetestmini 19.8 24.7 22.2 21.0 26.2
MathVision 15.4 12.6 16.3 17.5 18.4
Text-rich DocVQAtest 71.0 86.9 87.6 89.9 85.0
InfoVQAtest 40.0 59.5 59.1 65.4 58.3
ChartQAtest 59.6 71.4 57.1 73.5 76.5
TextVQAval 74.3 73.5 77.6 79.9 76.4
OCRVQAtestcore 54.4 40.2 67.6 68.7 64.0
VCRen easy 27.6 51.6 67.7 68.3 70.0
OCRBench 613 784 782 810 772
Avg Score 53.5 58.8 60.9 62.1 64.1
Table 5: Comprehensive Benchmark Comparisons of Aquila-VL-2B Model and State-of-the-art.
Appendix B Video Understanding
@@
Appendix C Data composition
We listed the sources, sizes, and types of all our data in Table 7.

### v2

@@ 3.3 Data Processing
After collecting all the data, we proceed with data processing. First, we standardize the format of data from various sources. Then, we remove duplicate Image-Text pairs and filter out images with high similarity based on their pHash values [75]. Besides, we use Qwen2-VL-2B [63] to calculate the loss for each sample and exclude the top 5% data with the highest loss, as high loss in well-trained multimodal models often indicates noisy data.
4 Training of Aquila-VL-2B
@@ 4.1 Model Architecture
To validate the effectiveness of Infinity-MM, we used it to train Aquila-VL-2B, a vision-language model with about 2 billion parameters. Aquila-VL-2B builds upon LLaVA-OneVision architecture[33], comprising a language tower, a vision tower, and a projector.
@@
  • Language Tower We choose Qwen-2.5-Instruct [6] as the language tower for its outstanding performance among open-source models.
@@
  • Vision Tower We utilize SigLIP [76], with approximately 400 million parameters, as the vision tower to extract visual features from input images and videos.
@@
  • Projector We utilize a two-layer MLP [41] with a GELU [27] activation to project visual features into the word embedding space.
Table 4: Comprehensive benchmark Comparisons of Aquila-VL-2B Model and other models. Results for models marked with * are sourced directly from the VLMEvalKit leaderboard, while other results are measured by us using default configurations. In all experimental results, higher values indicate better performance. The Aquila-VL-2B model, trained with Infinity-$M$, achieved SOTA performance on average among models of the same scale.
Model MiniMonkey*[28] PaliGemma*[8] DeepSeekVL*[44] H2OVL*[23] Phi3-V*[1] Vintern*[20] MiniCPM-V2[70] InternVL2[14] XinYuanVL[15] Qwen2VL[63] Aquila-VL
Pars (B) 2.2 2.9 2.0 2.1 4.2 3.7 2.8 2 2 2 2
Open Data
MMB-ENtest[42] 72.4 71.0 66.4 72.1 73.6 70.6 69.4 73.4 78.9 74.9 78.8
MMB-CNtest[42] 70.3 63.6 62.9 62.9 62.4 69.4 65.9 70.9 76.1 73.9 76.4
MMBV1.1[42] 68.9 65.6 63.8 64.8 65.2 66.6 65.2 69.7 75.4 72.7 75.2
MMT-Benchall[72] - - - - - - 54.5 53.3 57.2 54.8 58.2
RealWorldQA[67] 57.1 55.2 49.7 62.9 58.8 58.2 55.4 57.3 63.9 62.6 63.9
HalluB[24] 38.7 32.2 27.6 35.9 39.0 43.2 36.8 38.1 36.0 41.5 43.0
S-Bench2Plus[32] - 49.8 43.7 59.5 64.2 64.1 51.8 60.0 63.0 62.4 63.0
LLaVABench[39] 61.2 36.9 51.1 65.7 63.9 62.2 66.1 64.8 42.4 52.5 68.4
MMStar[13] 50.4 48.3 39.9 48.9 47.7 47.5 41.6 50.2 51.9 47.8 54.9
POPE[37] - 87.5 85.9 86.6 83.7 87.4 86.6 85.3 89.4 88.0 83.6
MMVet[73] 38.0 33.1 29.2 41.1 44.1 37.8 44.0 41.1 42.7 50.7 44.3
GeneralVQA MME[22] 1881.6 1686.1 1531.6 1767.5 1508.0 1782.6 1788.6 1863.0 1854.9 1890.0 1799.3
MMMUval[74] 35.0 34.9 33.8 36.3 46.1 46.7 39.6 34.9 43.6 41.7 47.4
S-QAtest[45] - 94.3 68.4 92.1 90.0 75.0 80.4 94.1 86.6 78.1 95.2
AI2Dtest[30] 74.8 68.3 51.5 70.9 78.4 69.1 64.8 74.4 74.2 74.6 75.0
MathVista[46] 45.2 28.7 29.8 56.5 44.6 43.4 39.0 45.0 47.1 47.9 59.0
MathVerse[77] - - - - - - 19.8 24.7 22.2 21.0 26.2
Knowledge&Math MathVision[62] - - - - - - 15.4 12.6 16.3 17.5 18.4
DocVQAtest[49] - - - - - - 71.0 86.9 87.6 89.9 85.0
InfoVQAtest[49] - - - - - - 40.0 59.5 59.1 65.4 58.3
ChartQAtest[48] - 33.7 47.4 59.4 81.8 68.3 59.6 71.4 57.1 73.5 76.5
TextVQAval[57] - 68.1 57.8 74.8 72.4 67.2 74.3 73.5 77.6 79.9 76.4
OCRVQAtest[50] - 57.8 58.1 70.3 61.9 56.8 54.4 40.2 67.6 68.7 64.0
VCRen easy[78] - - - - - - 27.6 51.6 67.7 68.3 70.0
Text-rich OCRBench [43] 804 614 413 778 637 618 613 784 782 810 772
Average - - - - - - 53.9 59.1 61.1 62.3 64.1
4.2 Training Details
We use the official codes of LLaVA-OneVision [^1] https://github.com/LLaVA-VL/LLaVA-NeXT/tree/main/scripts/train and the training setup is given in Table 3. Followling Li et al. [34], the training is divided into different stages, progressively increasing the difficulty, image resolution, and data quality:
  • Stage 2: We utilize general visual instruction data for further training to equip the model with fundamental capabilities for solving multimodal tasks. The data is divided into three subsets, and during each stage of training, the maximum visual resolution was progressively increased to enhance the model’s comprehension of visual information.
  • Stage 3: We employ selective visual instruction data for training and further increase the maximum resolution to improve performance.
  • Stage 4: We fine-tune the model using training data from GPT-4 and synthetic data. Experiments demonstrate that this part of data can further enhance model performance.
5 Evaluation
In this section, we first compared the performance of Aquila-VL-2B with similarly sized models across different benchmarks, showing that Aquila-VL-2B achieves SOTA performance. Next, we assess the impact of training with Infinity-MM versus other open-source datasets. Finally, we analyze the impact of our synthetic data on model performance and examine how performance varies with data scale.
5.1 Main Results
We assessed the visual capabilities of Aquila-VL-2B using a range of visual benchmarks provided by the VLMEvalKit [21]. Experimental results are shown in Table 4. Aquila-VL-2B demonstrates highly competitive performance at the same scale, achieving new state-of-the-art results. Specifically, we evaluated the capabilities of Aquila-VL-2B across three task categories.
General Visual Question Answering We conducted extensive evaluations across a diverse array of general visual question answering benchmarks: MMBench-ENtest (MMB-ENtest), MMBench-CNtest (MMB-CNtest), MMBench-1.1 (MMBV1.1) [42], MMT-Benchall [72], RealWorldQA [67], HallusionBench (HalluB) [24], SEEDBench2plus (S-Bench2plus) [32]), LLaVABench [39], MMStar [13],POPE [37] ,MMVet [73], and MME [22] datasets. Aquila-VL-2B demonstrates strong performance across these benchmarks, achieving or surpassing state-of-the-art results in most cases at the same scale. Specifically, Aquila-VL-2B achieves the best performance on MMB-CNtest, MMBV1.1, MMT-Benchall, RealWorldQA, LLaVABench, and MMSar. It also demonstrates top-tier results on MMB-ENtest, HalluB, and S-Bench2Plus. Additionally, it achieves competitive results on POPE, MMVet, and MME, with room for further improvement. These results strongly indicate Aquila-VL-2B’s remarkable capability in handling general visual question-answering tasks and validate the effectiveness of using the Infinity-MM dataset to enhance visual understanding.
Knowledge and Mathematical Reasoning We conducted experiments on MMMUval [74], ScienceQAtest (S-QAtest) [45], AI2Dtest [30], MathVista [46], MathVerse [77], and MathVision [62] to evaluate the model’s capabilities in knowledge and mathematical reasoning. Aquila-VL-2B achieves state-of-the-art performance on MMMUval, S-QAval, MathVista, MathVerse, and MathVision, while securing top-tier results on AI2Dtest. These outcomes underscore Aquila-VL-2B’s exceptional capabilities in visual knowledge comprehension and mathematical reasoning. The success can be attributed to Infinity-MM, which has collected over 2 million high-quality mathematical reasoning data points, further demonstrating the dataset’s quality and effectiveness.
Text Reading We assessed Aquila-VL-2B’s capabilities in text reading and diagram comprehension using DocVQAtest [49], InfoVQAtest [49], ChartQAtest [48], TextVQAval [57], OCRVQAtestcore [50] and VCRen easy [78] and OCRBench [43]. The experimental results indicate that Aquila-VL-2B’s text comprehension abilities are generally among the top tier, with potential for further improvement. We will continually supplement high-quality, text-rich data to further refine the Infinity-MM dataset.
Table 5: Performance comparison of models trained with Infinity-MM versus other open-source datasets. Compared to existing open-source datasets, Infinity-MM demonstrates a clear performance advantage.
@@ 5.1 Main Results
Model Params (B) Training Data Average MMBenchV1.1test MMStar MMMUval MathVistatestmini HallusionBench AI2Dtest OCRBench MMVet
LLaVA-OneVision-7B [33] 8 LLaVA-OneVision-SI 60.1 80.9 61.9 47.9 62.3 31.6 82.4 622 51.9
Cambrian-34B [60] 34 Cambrian-7M 58.3 77.8 54.2 50.4 50.3 41.6 79.5 591 53.2
Aquila-VL-2B 2 LLaVA-OneVision-SI 52.1 70.3 49.9 42.2 49.4 33.6 73.1 617 36.7
Infinity-$M$ 59.5 75.2 54.9 47.4 59.0 43.0 75.0 772 44.3
5.2 Infinity-MM vs. Other Open Datasets
To further validate the performance advantages of Infinity-MM and its importance for research in the open-source community, we compared Infinity-MM with several widely used, high-performance open-source datasets. We selected eight of the most representative datasets for testing, and the experimental results are shown in Table 5. Specifically, we first compare Aquila-VL-2B, trained on Infinity-MM, with LLaVA-OneVision-7B, trained on LlaVA-OneVision data, and Cambrian-34B, trained on Cambrian-10M. The results indicate that, despite its smaller scale relative to LLaVA-OneVision-7B and Cambrian-34B, Aquila-VL-2B outperforms Cambrian-34B and performs comparably to LLaVA-OneVision-7B. This underscores the critical role of training data in determining model performance. Furthermore, to better assess the relative impact of Infinity-MM versus LLaVA-OneVision data, we trained Aquila-VL-2B using each dataset independently. The results reveal that training on Infinity-MM yields superior performance, providing strong evidence of the high quality of the Infinity-MM dataset.
5.3 Effects of Synthetic Data
Table 6: Comparison of the model performance with and without synthetic data. The average performance of the model significantly declined without synthetic data, validating the effectiveness of synthetic data.
@@ 5.3 Effects of Synthetic Data
w/o Synthetic Data 57.25 75.09 54.53 45.56 57 35.89 75.03 766 38.3
@@
To assess the impact of synthetic data on model performance, we conducted an ablation study. In this experiment, we removed all synthetic data and trained the model using only the original GPT-generated data. The results, as shown in Table 6, reveal a significant decline in overall model performance after removing the synthetic data. This demonstrates that the synthetic data played a crucial role in enhancing the model’s performance, further validating the effectiveness of our approach in data augmentation and diversity.
@@ 5.4 Data Scaling
Figure 5: As the amount of instruction data used for training increased, the model’s average performance continued to improve, validating the effectiveness of scaling up instruction data.
To further analyze the impact of data size scaling on model performance, we conducted a detailed study on how model performance varies with the amount of training data. The results, shown in Figure 5, indicate a consistent improvement in performance as the training data increases. This trend clearly demonstrates that expanding the scale of instruction data has a significant positive effect on model performance. This observation indicates that incorporating more diverse instructional data enhances the Aquila-VL’s capability to tackle complex tasks. Consequently, expanding the volume of instructional data proves to be an effective strategy for boosting overall model performance.
@@ 6 Conclusion
In this work, we proposed the Infinity-MM multimodal instruction dataset, comprising tens of millions of samples to substantially increase data volume and enhance model effectiveness. Furthermore, we introduced a novel method for synthesizing instruction data based on open-source models, which enabled the generation of high-quality instruction data and further expanded the dataset. Finally, we trained the Aquila-VL-2B model on Infinity-MM, achieving state-of-the-art performance for models of a similar scale.
Limitations Despite the impressive performance of Infinitiy-MM, there is still room for enhancement in our work: (1) Due to limited computational resources, we have currently conducted experiments only on the 2B-scale model; (2) Due to time constraints, our dataset has not yet incorporated additional multi-image, video, and multilingual data. The future work will involve training on larger-scale models to further validate the quality of Infinity-MM. Additionally, we will continuously incorporate diverse high-quality data into Infinity-MM to support advancements in related research.
Appendix A Video Understanding
@@
Appendix B Data composition
For Image Caption Data, it consists entirely of General Instruction Data. But for Visual Instruction Data, due to its complex components, we further divide it into 3 parts, and visualize the data types of the three parts respectively, as shown in Figure 6. The composition of Comprehensive Data and Selective Data is General Instruction Data, OCR Data, Doc/Chart/Screen Data, Math/Reasoning Data and Text Instruction Data, for GPT4 & Synthetic Data, we add our Synthetic Data. Besides, we listed the sources, sizes, and types of all our data in Table 8.
(a) Comprehensive Data
(b) Selective Data
(c) GPT4 & Synthetic Data
Figure 6: Distribution of Visual Instruction Data